# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ 5.2 Wasserstein Bounds for the Signature Operation
${\mathbb{W}}^{2}_{2}\left(\sum_{i=1}^{M}\bar{\pi}^{i}\mathcal{N}(m_{i},\Sigma_{i}),\sum_{i=1}^{M}{\bar{\pi}^{{(i)}}}{\Delta}_{{{\mathcal{C}}}_{i}}\#\mathcal{N}(m_{i},\Sigma_{i})\right)\leq\sum_{i=1}^{M}{\bar{\pi}^{{(i)}}}{\mathbb{W}}^{2}_{2}\left(\mathcal{N}(m_{i},\Sigma_{i}),{\Delta}_{{{\mathcal{C}}}_{i}}\#\mathcal{N}(m_{i},\Sigma_{i})\right).$ (12)
That is, for a mixture of Gaussian distributions, the resulting 2-Wasserstein distance from a signature can be bounded by a weighted sum of the 2-Wasserstein distance between each of the Gaussian distributions in the mixture and their signatures. [^6] This trivial result is a special case of Proposition 24 in the Appendix on the Wasserstein distances between mixture distributions. In what follows, with an abuse of notation, for a Gaussian mixture distribution $p=\sum_{i=1}^{M}{\bar{\pi}^{{(i)}}}\mathcal{N}(m_{i},\Sigma_{i})$ we will refer to its signature as ${\Delta}_{\mathcal{C}}\#p:=\sum_{i=1}^{M}{\bar{\pi}^{{(i)}}}{\Delta}_{{{\mathcal{C}}}_{i}}\#\mathcal{N}(m_{i},\Sigma_{i})$. [^7] Note that an alternative approach would be to apply the same signature locations to all Gaussian distributions in the mixture. However, this would require integrating Gaussian distributions with a non-diagonal covariance over a hyper-rectangle, for which closed-form solutions such as those derived in Corollary 10 are not available, and this should hence be performed via numerical integration approaches.
Figure 4: Construction of the signature of a 2D Gaussian distribution (black dots) using the signature of a standard Gaussian distribution (blue dots) as a template. In the space induced by ${{T}}$, that is, the basis of the covariance matrix (orange arrows) of the Gaussian distribution, all dimensions of the Gaussian distribution are independent. Hence, we can take the cross-product of the signatures of each univariate Gaussian marginal in dimension $j$ of the transformed space with signature locations $\bar{{\mathcal{C}}}^{j}$ (green stripes) as a template. The signature in the original space is then obtained by taking the post image of the template under ${{T}}^{-1}$. The blue and black lines represent the edges of the Voronoi partition associated with the signature locations in the transformed and original space, respectively.
@@ 6 Stochastic Neural Networks as Gaussian Mixture Models
In this section, we detail and formalize our approach as illustrated in Figure 2 to iteratively approximate a SNN with a GMM. We first consider the case where ${\mathcal{X}}=\{{{x}}\}$, i.e., we approximate the distribution of a SNN over a single input point. The extension of our results for a finite set of input points will be considered in Section 6.3. Last, in Section 6.4, we prove that $q_{nn({\mathcal{X}})}$, i.e., the Gaussian mixture approximation resulting from our approach, converges uniformly to $p_{nn({\mathcal{X}})}$. That is, the error between $p_{nn({\mathcal{X}})}$ and $q_{nn({\mathcal{X}})}$ can be made arbitrarily small by increasing the number of signature points and GMM components.
@@ 6.1 Gaussian Mixture Model Approximation of a Neural Network
As shown in Eqn. (4), $p_{nn({{x}})}$, the output distribution of a SNN at input point ${{x}}$, can be represented as a composition of $K$ stochastic operations, one for each layer. To find a Gaussian mixture distribution $q_{nn({{x}})}$ that approximates $p_{nn({{x}})}$, our approach is based on iteratively approximating each of these operations with a GMM, as illustrated in Figure 2. Our approach can then be formalized as in Eqn (13) below for $k\in\{1,\ldots,K\}$:
$q_{nn(x),1}$ $={\mathbb{E}}_{{{{z}}}_{0}\sim\delta_{{x}}}[p_{L_{{{w}}_{0}}({{{z}}}_{0})}]$ (Initialization and Linear operation) (13a)
$d_{nn(x),k}$ $={\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({{x}}),k})$ (Compression and Signature operations) (13b)
$q_{nn(x),k+1}$ $={\mathbb{E}}_{{{{z}}}_{k}\sim d_{nn({{x}}),k}}[p_{L_{{{w}}_{k}}({\sigma}({{{z}}}_{k}))}]$ (Activation and Linear operations) (13c)
$=\sum_{i=1}^{|{\mathcal{C}}_{k}|}{\bar{{{\pi}}}^{{(i)}}}\mathcal{N}\big{(}\bar{{{C}}}_{i}{m}_{W_{k}}+{m}_{b_{k}},\bar{{{C}}}_{i}\Sigma_{W_{k}}\bar{{{C}}}_{i}^{T}+\Sigma_{b_{k}}\big{)}$
@@
where ${\mathcal{C}}_{k}\subset{\mathbb{R}}^{n_{k}}$ is the set of signature locations for layer $k$, $\bar{{{C}}}_{i}={{I}}_{n_{k+1}}\otimes{\sigma}({{c}}_{i}^{T})$ with ${{c}}_{i}\in{\mathcal{C}}_{k}$, and ${m}_{W_{k}},\Sigma_{W_{k}},{m}_{b_{k}},$ and $\Sigma_{b_{k}}$ are the means and the covariances of the weight matrix $W_{k}$ and bias vector $b_{k}$. ${R}_{M}:\mathcal{G}_{\infty}({\mathbb{R}}^{n_{k}})\rightarrow\mathcal{G}_{M}({\mathbb{R}}^{n_{k}})$ is the compression operation that compresses the Gaussian mixture distribution $q_{nn({{x}}),k}$ into a Gaussian Mixture distribution with at most $M$ components, thus limiting the complexity of the approximation. Details on how we perform operation $R_{M}$ using moment matching will be given in Section 7.1. For the rest of this section, ${R}_{M}(q_{nn({{x}}),k})$ can be simply considered as a mapping from the GMM $q_{nn({{x}}),k}$ into a GMM with a maximum size of $M$.
Eqn. (13) consists of the following steps: $q_{nn(x),1}$ is the distribution resulting from a linear combination of the input $x$ with the weights at the first layer, $d_{nn(x),k}$ is the result of a compression and signature operation of $q_{nn({{x}}),k}$, the output of the previous layer. The approximate output distribution $q_{nn(x),k+1}$ at each layer $k$ is obtained by marginalizing $p_{L_{{{w}}_{k}}({\sigma}({{{z}}}_{k}))}$ w.r.t. $d_{nn(x),k}$. We recall that for any fixed ${{{z}}}_{k}$, $p_{L_{{{w}}_{k}}({\sigma}({{{z}}}_{k}))}$ is Gaussian, as it is the distribution of a linear combination of vector ${\sigma}({{{z}}}_{k})$ with matrix $W_{k}$ and vector $b_{k}$, whose components are jointly Gaussian (and Gaussian distributions are closed w.r.t. linear combination). Consequently, as for all $k$, $d_{nn(x),k}$ is a discrete distribution, $q_{nn(x),k+1}$ is a Gaussian mixture distribution with a size equal to the support of $d_{nn(x),k}$.
@@
The approach described in Eqn (13) leads to a Gaussian mixture distribution $q_{nn({{x}})}$ under the assumption that the weight distribution $p_{{{w}}}$ is Gaussian. In the more general case, where $p_{{{w}}}$ is non-Gaussian, one can always recursively approximate $p_{L_{{{w}}_{k}}({{z}})}$ for each $k$ by a discrete distribution using the signature operation. That is, $p_{L_{{{w}}_{k}}({{z}})}$ in Eqn. (13c) is replaced with $q_{L_{{{w}}_{k}}({{z}})}=L_{{{w}}_{k}}({{z}})\#{\Delta}_{{\mathcal{C}}_{{{w}}_{k}}}\#p_{{{w}}_{k}}$, where ${\mathcal{C}}_{{{w}}_{k}}\subset{\mathbb{R}}^{(n_{k+1}+1)\times(n_{k}+1)}$ is the set of signature locations. Applying this additional operation, [^8] In Proposition 26 in Appendix A.3, it is shown how to formally account for the additional approximation error introduced by this additional signature operation. Eqn (13) will lead to a discrete distribution for $q_{nn({{x}})}$.
@@ 6.2 Wasserstein Distance Guarantees
Since $q_{nn(x)}$ is built as a composition of $K$ iterations of the stochastic operations in Eqn (13), to bound ${\mathbb{W}}_{2}(p_{nn(x)},q_{nn(x)})$ we need to quantify the error introduced by each of the steps in Eqn (13). A summary of how each of the operations occurring in Eqn (13) modifies the Wasserstein distance between two distributions is provided in Table 11 (details are given in Appendix A.3). These bounds are composed using interval arithmetic to bound ${\mathbb{W}}_{2}(p_{nn(x)},q_{nn(x)})$ in the following theorem.
@@
Let $p_{nn({{x}})}$ be the output distribution of a SNN with $K$ hidden layers for input ${{x}}\in{\mathbb{R}}^{n_{0}}$ and $q_{nn({{x}})}$ be a Gaussian mixture distribution built according to Eqn. (13). For $k\in\{1,\ldots,K\}$ iteratively define $\hat{{\mathbb{W}}}_{2}$ as
@@
$\hat{{\mathbb{W}}}_{2,k+1}=\hat{\mathcal{S}}_{2,{{w}},k}\left[{\mathcal{L}}_{{\sigma}}\hat{{\mathbb{W}}}_{2,k}+{\mathcal{L}}_{{\sigma}}\hat{{\mathbb{W}}}_{2,{R},k}+\hat{{\mathbb{W}}}_{2,{\Delta},k}\right],$ (14b)
@@
where $\hat{\mathcal{S}}_{2,{{w}},k}$, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ and $\hat{{\mathbb{W}}}_{2,{R},k}$ are such that
${\mathbb{E}}_{{{{W}}}_{k}\sim p_{w_{k}}}[\|{{{W}}}_{k}\|^{2}]^{\frac{1}{2}}$ $\leq\hat{\mathcal{S}}_{2,{{w}},k},$ (15)
${\mathbb{W}}_{2}({\sigma}\#{R}_{M}(q_{nn({{x}}),k}),{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({{x}}),k}))$ $\leq\hat{{\mathbb{W}}}_{2,{\Delta},k},$ (16)
${\mathbb{W}}_{2}(q_{nn({{x}}),k},{R}_{M}(q_{nn({{x}}),k}))$ $\leq\hat{{\mathbb{W}}}_{2,{R},k}.$ (17)
Then, it holds that
${\mathbb{W}}_{2}(p_{nn(x)},q_{nn(x)})\leq\hat{{\mathbb{W}}}_{2}(p_{nn(x)},q_{nn(x)})\coloneqq\hat{{\mathbb{W}}}_{2}.$ (18)
The proof of Theorem 13 is reported in Appendix A.3 and is based on the triangular inequality property of the 2-Wasserstein distance, which allows us to bound ${\mathbb{W}}_{2}(p_{nn({{x}})},q_{nn({{x}})})$ iteratively over the hidden layers. Note that Theorem 13 depends on quantities $\hat{\mathcal{S}}_{2,{{w}},k}$, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ and $\hat{{\mathbb{W}}}_{2,{R},k}$, which can be computed as follows. First, $\hat{\mathcal{S}}_{2,{{w}},k}$ is a bound on the expected spectral norm of the weight matrices of the $k$-th layer, which can be obtained according to the following Lemma.
Lemma 14
For distribution $p_{{W}}\in{\mathcal{P}}({\mathbb{R}}^{n\times m})$, we have that
${\mathbb{E}}_{{{W}}\sim p_{{W}}}[\|{{W}}\|^{2}]^{\frac{1}{2}}$ $\leq\Big{(}\sum_{i,j}{\mathbb{E}}_{\bar{{{W}}}\sim\tau_{{M}}\#p_{{W}}}[({\bar{{{W}}}^{{(i,j)}}})^{2}]\Big{)}^{\frac{1}{2}}+\|{{M}}\|$ (19)
where $\tau_{{M}}:{\mathbb{R}}^{n\times m}\rightarrow{\mathbb{R}}^{n\times m}$ is a translation defined by $\tau_{{M}}({{W}})={{W}}+{{M}}$ and ${{M}}={\mathbb{E}}_{{{W}}\sim p_{{W}}}[{{W}}]$, i.e., $\tau_{{M}}\#p_{{W}}$ is a zero mean distribution.
The proof of Lemma 14 relies on the Minkowski inequality and the fact that the spectral norm is upper bounded by the Frobenius norm, as explained in Appendix A.3. Furthermore, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ can be computed as in Eqn. (12) using Corollaries 10 and 27. Lastly, $\hat{{\mathbb{W}}}_{2,{R},k}$ bounds the distance between Gaussian mixture distributions, which can be efficiently bounded using the ${\mathbb{MW}}_{2}$ distance introduced in Delon and Desolneux (2020) and formally defined below.
Definition 15 (${\mathbb{MW}}_{2}$ Distance)
Let $p=\sum_{i=1}^{N_{p}}{{{\pi}}_{p}^{{(i)}}}p_{i}\in\mathcal{G}_{N_{1}}({\mathbb{R}}^{n})$ and $q=\sum_{j=1}^{N_{q}}{{{\pi}}_{q}^{{(j)}}}q_{j}\in\mathcal{G}_{N_{1}}({\mathbb{R}}^{n})$ be two Gaussian Mixture distributions. The ${\mathbb{MW}}_{2}$-distance is defined as
${\mathbb{MW}}_{2}(p,q)\coloneqq\left(\inf_{\gamma\in\Gamma(p,q)\cap\mathcal{G}_{<\infty}({\mathbb{R}}^{n}\times{\mathbb{R}}^{n})}{\mathbb{E}}_{({{x}},{{z}})\sim\gamma}[\|{{x}}-{{z}}\|^{2}]\right)^{\frac{1}{2}}=\bigg{(}\min_{{{\pi}}\in\Gamma({{\pi}}_{p},{{\pi}}_{q})}\sum_{i}^{N_{p}}\sum_{j}^{N_{q}}{{{\pi}}^{{(i,j)}}}{\mathbb{W}}_{2}^{2}(p_{i},q_{j})\bigg{)}^{\frac{1}{2}}.$ (20)
The ${\mathbb{MW}}_{2}$ distance is a Wasserstein-type distance that restricts the set of possible couplings to Gaussian mixtures. Consequently, it can be formulated as a finite discrete linear program with $N_{p}\cdot N_{q}$ optimization variables and coefficients being the Wasserstein distance between the mixture’s Gaussian components, which have closed-form expressions (Givens and Shortt, 1984).999 For $p_{1}=\mathcal{N}({m}_{1},\Sigma_{1})\in\mathcal{G}({\mathbb{R}}^{n})$ and $p_{2}=\mathcal{N}({m}_{2},\Sigma_{2})\in\mathcal{G}({\mathbb{R}}^{n})$, it holds that ${\mathbb{W}}_{2}^{2}(p_{1},p_{2})=\|{m}_{1}-{m}_{2}\|^{2}+\mathrm{trace}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{1}^{\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{\frac{1}{2}})^{\frac{1}{2}})$.
Operation Wasserstein Bounds Ref.
Linear
${\mathbb{W}}_{2}({\mathbb{E}}_{{{x}}\sim p}[p_{L_{{{w}}}({{x}})}],{\mathbb{E}}_{{{z}}\sim q}[p_{L_{{{w}}}({{z}})}])$ $\leq{\mathbb{E}}_{({{W}},{{b}})\sim p_{w}}[\|{{W}}\|^{2}]^{\frac{1}{2}}{\mathbb{W}}_{2}(p,q)$ Prop. 25
Activation
${\mathbb{W}}^{2}_{2}({\sigma}\#p,{\sigma}\#{\Delta}_{\mathcal{C}}\#p)$ $\leq\sum_{i=1}^{|{\mathcal{C}}|}{\mathcal{L}}^{2}_{{\sigma}\mid{\mathcal{R}}_{i}}{\mathbb{W}}_{2\mid{\mathcal{R}}_{i}}^{2}$ Corol. 27
${\mathbb{W}}_{2}({\sigma}\#p,{\sigma}\#q)$ $\leq{\mathcal{L}}_{{\sigma}}{\mathbb{W}}_{2}(p,q)$
Signature/ Compression $h\in\{{\Delta}_{\mathcal{C}},{R}_{R}\}$
${\mathbb{W}}_{2}(p,h\#q)$ $\leq{\mathbb{W}}_{2}(p,q))+{\mathbb{W}}_{2}(q,h\#q)$ Prop. 28

### v2

@@ 5.3 Convergence Rates for Signature Approximations of Gaussian Mixtures
${\mathbb{W}}_{2}^{2}\left(p,\bm{{\Delta}}_{\{{\mathcal{C}}_{i}\}_{i=1}^{M}}\#p\right)\leq\sum_{i=1}^{M}{\tilde{\pi}^{{(i)}}}\sum_{j=1}^{n}{{{\lambda}}_{i}^{{(j)}}}\left(\frac{4\log N_{i,j}}{N_{i,j}^{2}}+r(N_{i,j})\right)$ (18)
where $\lambda$}} is the vector of eigenvalues of $\Sigma_{i}$.
Corollary 15 follows from Eqn (14), which decomposes the mixture error into a mixture-weighted sum of component errors. For each component we diagonalize $\Sigma_{i}$, apply Corollary 11 to separate dimensions, and then invoke Proposition 13 on each one-dimensional uniform grid of points ${\mathcal{C}}_{i}^{j}$, yielding the closed-form expression for the error bound after summing over $i$ and $j$.
In Figure 5, we examine the convergence rate of the approximation error for 2D Gaussian mixtures decays as the number of signature locations increases, comparing three constructions: uniform Cartesian product grids (Corollary 15), adaptive (non‑uniform) grids, and an empirical proxy for the optimal signature locations. Figure 5(a) confirms that adaptive grids substantially reduce the ${\mathbb{W}}_{2}$ error relative to the uniform case (see Remark 14). Figure 5(b) further shows that adaptive placement already captures most of the attainable improvement, with only modest additional gains from unconstrained (cluster-based) locations for the distinct-mode mixtures considered. Appendix F provides further discussion and additional experiments on the conservatism of the component‑wise signature bound in Eqn. (14).
@@ 6 Stochastic Neural Networks as Gaussian Mixture Models
In this section, we detail and formalize our approach as illustrated in Figure 2 to iteratively approximate a SNN with a GMM at a finite set of input points $\mathcal{X}$}. We first consider the case where ${\mathcal{X}}=\{{{x}}\}$, i.e., we approximate the distribution of a SNN over a single input point. The extension to finite set of input points is considered in Section 6.3. Last, in Section 6.4, we prove that $q_{nn({\mathcal{X}})}$, i.e., the Gaussian mixture approximation resulting from our approach, converges to $p_{nn({\mathcal{X}})}$ in Wasserstein distance. That is, the error between $p_{nn({\mathcal{X}})}$ and $q_{nn({\mathcal{X}})}$ can be made arbitrarily small by increasing the number of signature locations and GMM components.
@@ 6.1 Gaussian Mixture Model Approximation of a Neural Network
As shown in Eqn (4), $p_{nn({{x}})}$, the output distribution of a SNN at input point x{{x}}, can be represented as a composition of $K$ stochastic operations, one for each layer. To find a Gaussian mixture distribution $q_{nn({{x}})}$ that approximates $p_{nn({{x}})}$, our approach is based on iteratively approximating each of these operations with a GMM, as illustrated in Figure 2. Our approach can then be formalized as in Eqn (19) below for $k\in\{1,\ldots,K\}$:
$q_{nn(x),1}$ $={\mathbb{E}}_{{{{z}}}_{0}\sim\delta_{{x}}}[{L^{{{w}}_{0}}({{{z}}}_{0})\#p_{{{w}}_{0}}}]$ (Initialization and Affine operation) (19a)
$d_{nn(x),k}$ $={\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\#{G}_{{M_{k}}}(q_{nn({{x}}),k})$ (Compression and Signature operations) (19b)
$q_{nn(x),k+1}$ $={\mathbb{E}}_{{{{z}}}_{k}\sim d_{nn({{x}}),k}}[{L^{{{w}}_{k}}({\sigma}({{{z}}}_{k}))\#p_{{{w}}_{k}}}]$ (Activation and Affine operations) (19c)
@@
where $\bm{{\mathcal{C}}}_{k}=\{{\mathcal{C}}_{k,i}\subset{\mathbb{R}}^{n_{k}}\}_{i=1}^{M_{k}}$ denotes the signature locations for layer $k$. ${G}_{M_{k}}:\mathcal{G}_{\infty}({\mathbb{R}}^{n_{k}})\rightarrow\mathcal{G}_{{M_{k}}}({\mathbb{R}}^{n_{k}})$ is a compression operation that compresses the Gaussian mixture distribution $q_{nn({{x}}),k}$ into a Gaussian Mixture distribution with at most $M_{k}$ components, thus limiting the approximation’s complexity. We show how operation ${G}_{{M_{k}}}$ is performed using moment matching in Section 7.1.
Eqn (19) consists of the following steps: $q_{nn(x),1}$ is the distribution resulting from an affine combination of the input $x$ with the parameters at the first layer, $d_{nn(x),k}$ is the result of a compression and signature operation of $q_{nn({{x}}),k}$, the output of the previous layer. The approximate output distribution $q_{nn(x),k+1}$ at each layer $k$ is obtained by marginalizing $L^{{{w}}_{k}}({\sigma}({{{z}}}_{k}))\#p_{{{w}}_{k}}$ w.r.t. $d_{nn(x),k}$. We recall that for any fixed ${{{z}}}_{k}$, and Gaussian distribution $p_{{{w}}_{k}}$, $L^{{{w}}_{k}}({\sigma}({{{z}}}_{k}))\#p_{{{w}}_{k}}$ is Gaussian, as Gaussian distributions are closed w.r.t. affine combination. Consequently, as for all $k$, $d_{nn(x),k}$ is a discrete distribution, $q_{nn(x),k+1}$ is a Gaussian mixture distribution with a size equal to the support of $d_{nn(x),k}$.
Example 1
Consider the case where $L^{{{w}}_{k}}$ is a fully-connected layer with weight matrix ${{W}}\in{\mathbb{R}}^{n_{k+1}\times n_{k}}$ and bias vector ${{b}}\in{\mathbb{R}}^{n_{k+1}}$, i.e., $L^{{{w}}_{k}}({{z}})={{W}}{{z}}+{{b}}$. Assume ${\mathrm{vec}}({{W}})$ and b{{b}} are distributed according to $\mathcal{N}({\mathrm{vec}}({{M}}),\Sigma)$ and $\mathcal{N}({{m}}_{{{b}}},\Sigma_{{{b}}})$, respectively, where ${\mathrm{vec}}(\cdot)$ stacks the columns of a matrix into a single vector. Then, the approximate output distribution at the $(k+1)$-th layer, $q_{nn(x),k+1}$, as defined in Eqn (19c) for $M_{k}=1$, signature locations ${\mathcal{C}}_{k,1}=\{{{c}}_{i}\}_{i=1}^{N_{k}}$ and partition $\{{\mathcal{R}}_{i}\}_{i=1}^{N_{k}}$ from Eqn (7), can be written as
$q_{nn(x),k+1}=\sum_{i=1}^{N_{k}}\mathbb{P}_{{{z}}\sim{G}_{1}(q_{nn(x),k})}[{{z}}\in{\mathcal{R}}_{i}]\cdot\mathcal{N}\Bigg{(}{{m}}_{{{b}}}+{{M}}{\sigma}({{c}}_{i}),\Sigma_{{{b}}}+\sum_{j,l=1}^{n_{k+1}}{{\sigma}({{c}}_{i})^{{(j)}}}{{\sigma}({{c}}_{i})^{{(l)}}}\Sigma_{jl}\Bigg{)}$ (20)
where $\Sigma_{jl}$ denotes the $(j,l)$-th block of the covariance matrix $\Sigma$, each block of size $n_{k}\times n_{k}$, i.e., the covariance between the $j$-th and $l$-th columns of W{{W}}.
Note that the approximation scheme in Eqn (19) allows for an arbitrary choice of signature locations ${\mathcal{C}}_{k,i}$ for the $i$-th component at each layer $k$. To obtain a closed-form expression for $q_{nn({{x}})}$, the locations $\bm{{\mathcal{C}}}_{k}$ should be selected as discussed in Remark 7; that is, so that $\mathbb{P}_{{{z}}\sim{G}_{1}(q_{nn(x),k})}[{{z}}\in{\mathcal{R}}]$ in Example 1 is analytically tractable for all regions $\mathcal{R}$}.
@@
The approach described in Eqn (19) leads to a Gaussian mixture distribution $q_{nn({{x}})}$ under the assumption that the parameter distributions $p_{{{w}}_{k}}$ are Gaussian. In the more general case, where $p_{{{w}}_{k}}$ is non-Gaussian, one can always recursively approximate $L^{{{w}}_{k}}({{z}})\#p_{{{w}}_{k}}$ for each $k$ by a discrete distribution using the signature operation. That is, $L^{{{w}}_{k}}({{z}})\#p_{{{w}}_{k}}$ in Eqn (19c) is replaced with $L^{{{w}}_{k}}({{z}})\#{\Delta}_{{\mathcal{C}}_{{{w}}_{k}}}\#p_{{{w}}_{k}}$, where ${\mathcal{C}}_{{{w}}_{k}}$ is the set of signature locations. Applying this additional operation, Eqn (19) will lead to a discrete distribution for $q_{nn({{x}})}$.
@@ 6.2 Wasserstein Distance Guarantees
Since $q_{nn(x)}$ is built as a composition of $K$ iterations of the stochastic operations in Eqn (19), to bound ${\mathbb{W}}_{2}\left(p_{nn(x)},q_{nn(x)}\right)$ we need to formally quantify the error introduced by each of the steps in Eqn (19). A summary of how each of the operations occurring in Eqn (19) modifies the Wasserstein distance between two distributions is provided in Table 13 (details are given in Appendix B.2). These bounds are then composed using interval arithmetic to bound ${\mathbb{W}}_{2}\left(p_{nn(x)},q_{nn(x)}\right)$ in the following theorem.
@@
Let $p_{nn({{x}})}$ be the output distribution of a SNN with $K$ hidden layers for input ${{x}}\in{\mathbb{R}}^{n_{0}}$ and $q_{nn({{x}})}$ be a Gaussian mixture distribution built according to Eqn (19). Iteratively define $\hat{{\mathbb{W}}}_{2}$ as
@@
$\hat{{\mathbb{W}}}_{2,k+1}={\mathcal{L}}_{k}\left[\hat{{\mathbb{W}}}_{2,k}+{\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right$.
$\hskip 142.26378pt\left.+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right],\qquad\forall k\in\{1,\ldots,K\},$ (21b)
@@
where $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}}}$ is defined in Eqn (14), ${\mathbb{MW}}_{2}$ as defined in Eqn (2), and
${\mathcal{L}}_{k}\coloneqq{\mathcal{L}}_{\sigma}{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}]^{\frac{1}{2}},$ (22)
with ${\mathcal{L}}_{\sigma}$ and ${\mathcal{L}}_{L^{{{w}}_{k}}}$ the Lipschitz constants of the activation function $\sigma$} and affine operation $L^{{{w}}_{k}}$, respectively. Then, it holds that
${\mathbb{W}}_{2}\left(p_{nn(x)},q_{nn(x)}\right)\leq\hat{{\mathbb{W}}}_{2}(p_{nn(x)},q_{nn(x)})\coloneqq\hat{{\mathbb{W}}}_{2}.$ (23)
The proof of Theorem 17 is reported in Appendix B.2 and is based on the triangle inequality property of the 2-Wasserstein distance, which allows us to bound ${\mathbb{W}}_{2}\left(p_{nn({{x}})},q_{nn({{x}})}\right)$ iteratively over the hidden layers. Note that Theorem 17 depends on the quantities ${\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}]^{1/2}$, $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{M_{k}}(q_{nn({{x}}),k})\right)$, and ${\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{M_{k}}(q_{nn({{x}}),k})\right)$, which can be computed as follows. First, ${\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}]^{1/2}$}, is the expectation of any Lipschitz constant ${\mathcal{L}}_{L^{{w}}}$ (any constant satisfying the Lipschitz inequality, not necessarily the minimal one) of the affine operation $L^{{{w}}_{k}}$, for which closed-form expressions for both feedforward and convolutional layers are given in Appendix C. Second, $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{M_{k}}(q_{nn({{x}}),k})\right)$ can be computed using Corollary 11, when the sets of signature locations satisfy the grid-regularity constraint in Eqn (12). Finally, ${\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)$ is obtained by computing the ${\mathbb{MW}}_{2}$-distance between Gaussian mixtures distributions (see Definition 2), which requires solving a finite discrete linear program with $M_{k}$ times the number of components of $q_{nn({\mathcal{X}}),k}$ optimization variables and coefficient given by the pairwise Wasserstein distances between the mixture’s Gaussian components.
Remark 18
If $p_{{{w}}_{k}}$ is non-Gaussian, the additional approximation step introduced in Remark 16 is accounted for by an extra application of the triangle inequality, which introduces an additional error term ${\mathbb{W}}_{2}(p_{{{w}}_{k}},{\Delta}_{{\mathcal{C}}_{{{w}}_{k}}}\#p_{{{w}}_{k}})$ in Eqn (21b). While the conditional expectations in Proposition 9 may no longer admit closed-form expressions in this case, they can always be conservatively estimated using approximate numerical integration.
Remark 19
In Theorem 17, instead of bounding the error from the signature approximation of ${G}_{{M_{k}}}(q_{nn({{x}}),k})$ in Eqn (21b), propagated through the activation function $\sigma$}, using the global Lipschitz constant ${\mathcal{L}}_{\sigma}$ of $\sigma$}, one can alternatively use a bound based on the local Lipschitz constant ${\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}$ w.r.t. region ${\mathcal{R}}_{i}$, as reported in Table 13. Specifically, one can replace ${\mathcal{L}}_{\sigma}\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({{x}}),k})\right)$ with
$\left(\sum_{i=1}^{N}{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}^{2}{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim{G}_{{M_{k}}}(q_{nn({{x}}),k})}[\|{{z}}-{{c}}_{i}\|^{\rho}\mid{{z}}\in{\mathcal{R}}_{i}]\right)^{1/2}$
in Eqn (21b). While generally tighter, this bound requires explicit computation of the local Lipschitz constant and the expectation w.r.t each region ${\mathcal{R}}_{i}$, which limits the efficiency of the grid structure ${\mathcal{C}}_{i}$ used in Corollary 11.
Operation Wasserstein Bounds
Affine ${\mathbb{W}}_{2}^{2}\left({\mathbb{E}}_{\tilde{{z}}\sim p}\left[L^{{{w}}}(\tilde{{z}})\#p_{{{w}}_{k}}\right],{\mathbb{E}}_{{{z}}\sim q}\left[L^{{{w}}}({{z}})\#p_{{{w}}}\right]\right)\leq{\mathbb{E}}_{{{w}}\sim p_{{{w}}}}\left[{\mathcal{L}}_{L^{{{w}}}}^{2}\right]{\mathbb{W}}_{2}^{2}(p,q)$
Activation ${\mathbb{W}}_{2}({\sigma}\#p,{\sigma}\#q)\leq{\mathcal{L}}_{{\sigma}}{\mathbb{W}}_{2}(p,q)$
${\mathbb{W}}^{2}_{2}({\sigma}\#p,{\sigma}\#{\Delta}_{\mathcal{C}}\#p)\leq{\sum_{i=1}^{N}{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}^{2}{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-{{c}}_{i}\|^{2}\mid{{z}}\in{\mathcal{R}}_{i}]}$
Signature/
Compression
${\mathbb{W}}_{2}(p,h\#q)\leq{\mathbb{W}}_{2}(p,q)+{\mathbb{W}}_{2}(q,h\#q)$, where $h\in\{{\Delta}_{\mathcal{C}},{G}_{M}\}$